### v1

# Optimal Quoting under Adverse Selection and Price Reading

Abstract
Over the past decade, many dealers have implemented algorithmic models to automatically respond to RFQs and manage flows originating from their electronic platforms. In parallel, building on the foundational work of Ho and Stoll, and later Avellaneda and Stoikov, the academic literature on market making has expanded to address trade size distributions, client tiering, complex price dynamics, alpha signals, and the internalization versus externalization dilemma in markets with dealer-to-client and interdealer-broker segments. In this paper, we tackle two critical dimensions: adverse selection, arising from the presence of informed traders, and price reading, whereby the market maker’s own quotes inadvertently reveal the direction of their inventory. These risks are well known to practitioners, who routinely face informed flows and algorithms capable of extracting signals from quoting behavior. Yet they have received limited attention in the quantitative finance literature, beyond stylized toy models with limited actionability. Extending the existing literature, we propose a tractable and implementable framework that enables market makers to adjust their quotes with greater awareness of informational risk.
Key words: Market Making, Stochastic Optimal Control, Adverse Selection, Price Reading, Informational Risk.
1 Introduction
In modern dealer markets, liquidity provision is increasingly handled by algorithmic systems that autonomously respond to clients’ RFQs and manage trade flows across multiple platforms. The academic literature has accompanied this technological shift with increasingly refined models. Building on the seminal contributions of Ho and Stoll [15] and Avellaneda and Stoikov [1], researchers have developed several models based on optimal control to deliver bid and ask price ladders to various types of clients, while managing inventory risk through quote skewing (internalization) and, when possible, hedging on trading venues (externalization); see [3, 7, 9, 13, 14].
Some models feature complex price dynamics [17], others incorporate alpha signals; see [8]. Several focus on multi-asset market making under correlation or cointegration assumptions; see [4, 5]. On the numerical side, substantial progress has been made, and reference methods now enable the approximation of optimal quotes in many contexts without suffering from the curse of dimensionality; see [6].
Yet, for all their sophistication, most existing market making models overlook two key challenges that are central to real-world practice: adverse selection and information leakage through price reading. Adverse selection arises when the market maker trades against participants whose order flow reflects superior information about future price dynamics. In other words, the market maker faces the risk of being hit at the bid (respectively ask) just before the market goes down (respectively up) – the so-called winner’s curse. Information leakage, as captured here by price reading, occurs when the market maker’s own quotes indirectly reveal information about their inventory or at least its sign – allowing certain traders, often informally referred to as skew sniffers (see [12]), to anticipate future market flows and subsequently build potentially profitable strategies.
While these risks are well known to practitioners (see [2]), they are rarely addressed in the quantitative finance literature. When they are, it is typically through highly stylized and very often static economic models, offering little guidance for implementation; see for instance [11]. [^5] Among the exceptions, [10] is an interesting paper which can be helpful to address the question of toxic flow, intimately related to our question of adverse selection. In particular, the interplay between informational risk and real-time quoting decisions remains insufficiently studied.
The goal of this paper is to help fill that gap. We introduce a tractable and implementable optimal market making model à la Cartea-Jaimungal, in infinite horizon, that accounts for both adverse selection and price reading effects. To avoid introducing a lot of assumptions, we rely on Taylor expansions and focus on the first-order adjustments in optimal quotes induced by each effect. The resulting equations provide practical insight into how market makers can adapt their behavior to flow toxicity and mitigate the risk of information leakage, while remaining competitive.
The remainder of the paper is organized as follows. Section 2 introduces the model. In Section 3, we derive the first-order effects of adverse selection and price reading on both the value function and the optimal quotes. Section 4 discusses informational effects in the context of a quadratic approximation of the Hamiltonian functions. Finally, Section 5 illustrates our findings in the case of simple functional forms.
2 Model Setup
Throughout this paper, we work on a filtered probability space $\left(\Omega,\mathcal{F},\mathbb{P};\mathbb{F}=\left(\mathcal{F}_{t}\right)_{t\geq 0}\right)$ satisfying the usual conditions. We assume that this space is rich enough to support all the stochastic processes introduced in the model.
2.1 Modeling Framework and Notation
We consider a market maker responsible for quoting bid and ask prices on a single asset. The market maker continuously sets quotes and can stream different pricing ladders to different categories of clients. Specifically, clients are divided into $N$ tiers, indexed by the set $\mathcal{N}:=\{1,\ldots,N\}$.
Transaction sizes are assumed to be discrete, taking values $0<\Delta^{1}<\ldots<\Delta^{K}$ for some $K\geq 1$, with index set $\mathcal{K}:=\{1,\ldots,K\}$. We assume $\Delta^{k}\in\Delta^{1}\mathbb{N}$ for all $k\in\mathcal{K}$, i.e., each size is an integer multiple of the smallest one.
For each tier $n\in\mathcal{N}$ and size index $k\in\mathcal{K}$, the bid and ask prices quoted at time $t$ are modeled by stochastic processes $\big{(}S^{n,k,b}_{t}\big{)}_{t\geq 0}$ and $\big{(}S^{n,k,a}_{t}\big{)}_{t\geq 0}$, respectively. These quotes determine the distribution of trade arrivals, which will be formalized later.
We introduce a reference price process $\big{(}S_{t}\big{)}_{t\geq 0}$, representing a composite price computed by the market maker based on public and private information. It evolves under both exogenous volatility and endogenous reactions to trading activity and quote asymmetry. Its dynamics is given by:
$\displaystyle dS_{t}=\$$\displaystyle\sigma\,dB_{t}-\sum_{\begin{subarray}{c}n\in\mathcal{N}\\ k\in\mathcal{K}\end{subarray}}\left(\tilde{\zeta}^{n,k}\left(S_{t}-S^{n,k,b}_{t}\right)\,dN^{n,k,b}_{t}-\tilde{\zeta}^{n,k}\left(S^{n,k,a}_{t}-S_{t}\right)\,dN^{n,k,a}_{t}\right)$ (1)
$\displaystyle+\sum_{n\in\mathcal{N}}\tilde{J}^{n}\left(\sum_{k\in\mathcal{K}}w^{n,k}\left(S^{n,k,a}_{t}-2S_{t}+S^{n,k,b}_{t}\right)\right)\,dt,\qquad S_{0}\text{ given,}$
where $\sigma>0$ is a volatility parameter and $\big{(}B_{t}\big{)}_{t\geq 0}$ is a standard Brownian motion adapted to the filtration $\left(\mathcal{F}_{t}\right)_{t\geq 0}$. The point processes $\big{(}N^{n,k,b}_{t}\big{)}_{t\geq 0}$ and $\big{(}N^{n,k,a}_{t}\big{)}_{t\geq 0}$ represent the number of trades of size $\Delta^{k}$ at the bid and ask with clients from tier $n$, respectively.
The model includes two key features that govern the endogenous reaction of the reference price to trade flow and quoting behavior:
  • Adverse selection. For each $(n,k)\in\mathcal{N}\times\mathcal{K}$, the function $\tilde{\zeta}^{n,k}:\mathbb{R}\to\mathbb{R}_{+}$ is non-decreasing, and satisfies $\lim_{\delta\to-\infty}\tilde{\zeta}^{n,k}(\delta)=0.$ This function captures the impact of informed trading on prices. It is tier-specific (and possibly size-specific), allowing the market maker to account for varying levels of expected informativeness across clients (and transaction sizes). In particular, when $\tilde{\zeta}^{n,k}$ is a rapidly decreasing function, executions at prices that deviate significantly from the reference price in the “wrong” direction – that is, at prices seemingly unfavorable to the client from the market maker’s perspective – can have a strong impact on the reference price.
  • Price reading. For each $n\in\mathcal{N}$, the function $\tilde{J}^{n}:\mathbb{R}\to\mathbb{R}$ is non-decreasing and satisfies $\tilde{J}^{n}(0)=0$. It models the way the reference price reacts to asymmetry in the quotes streamed to tier $n$. This asymmetry, or skew, is computed as a weighted average over the quote skewness for each trade size $\Delta^{k}$, where the inner term $S^{n,k,a}_{t}-2S_{t}+S^{n,k,b}_{t}$ captures the relative placement of bid and ask quotes with respect to the reference price. The weights $w^{n,\cdot}$ control the contribution of each size in the price ladder to the total skew associated with the quotes streamed to clients of tier $n$. This mechanism captures the phenomenon commonly referred to as price reading or skew sniffing, in which certain market participants infer the direction – or evolution – of the market maker’s inventory from the quote ladder structure and adjust their trading behavior accordingly, generating a feedback loop that moves the reference price.
Remark 1.
We define price reading effect via weighted average skew which is basically a supply-demand imbalance which universally appears in all short-term price prediction models and is sufficiently general (see [16] and [18]). However, other functional forms can be considered, particularly in the context of multi-dealer platforms where only the competitive side can be of importance.
The market maker’s inventory process $\left(q_{t}\right)_{t\geq 0}$ evolves according to
$dq_{t}=\sum_{\begin{subarray}{c}n\in\mathcal{N}\\ k\in\mathcal{K}\end{subarray}}\Delta^{k}\left(dN^{n,k,b}_{t}-dN^{n,k,a}_{t}\right),\quad q_{0}\text{ given}.$ (2)
The processes $\left(N^{n,k,b}_{t}\right)_{t\geq 0}$ and $\left(N^{n,k,a}_{t}\right)_{t\geq 0}$ have respective intensities $\left(\lambda^{n,k,b}_{t}\right)_{t\geq 0}$ and $\left(\lambda^{n,k,a}_{t}\right)_{t\geq 0}$, defined by
$\lambda^{n,k,b}_{t}=\Lambda^{n,k,b}\left(S_{t}-S^{n,k,b}_{t}\right)\quad\text{and}\quad\lambda^{n,k,a}_{t}=\Lambda^{n,k,a}\left(S^{n,k,a}_{t}-S_{t}\right).$ (3)
Here, $\Lambda^{n,k,b}$ and $\Lambda^{n,k,a}$ are smooth and decreasing functions, satisfying standard assumptions in market making models à la Avellaneda-Stoikov or Cartea-Jaimungal (see [13]); in particular,
$c^{n,k,b/a}(\delta):=2-\frac{\Lambda^{n,k,b}(\delta)\Lambda^{n,k,b^{\prime\prime}}(\delta)}{\Lambda^{n,k,b^{\prime}}(\delta)^{2}}$
verifies $\inf_{\delta}c^{n,k,b/a}(\delta)>0$.
Remark 2.
Although the reference price – whose dynamics is given by Eq. (1) – is primarily driven by its exogenous Brownian component, it also depends on the market maker’s clients’ flows and own quoting behavior. As such, it is a price specific to the market maker under consideration. In light of this, it may seem questionable at first glance to define the skew terms in Eq. (1) relative to this very reference price, and to let the intensities in Eq. (3) – which model the agents’ demand and supply behavior – also depend on the market maker’s own reference price. This modeling choice is, in fact, a conservative one: we assume that the market maker never benefits from any informational advantage, even though they observe clients’ flows. While restrictive, this assumption appears natural in the context of a paper focused on the informational risk faced by market makers. It is recognized that adverse selection can originate from the market maker not knowing the true reference price at the right time due to natural market fragmentation and the ensuing latency.
The market maker’s cash process $\left(X_{t}\right)_{t\geq 0}$ evolves according to:
$dX_{t}$ $=\sum_{\begin{subarray}{c}n\in\mathcal{N}\\ k\in\mathcal{K}\end{subarray}}\Delta^{k}\left(-S^{n,k,b}_{t}\,dN^{n,k,b}_{t}+S^{n,k,a}_{t}\,dN^{n,k,a}_{t}\right)$ (4)
$=\sum_{\begin{subarray}{c}n\in\mathcal{N}\\ k\in\mathcal{K}\end{subarray}}\Delta^{k}\left(-\left(S_{t}-\delta^{n,k,b}_{t}\right)\,dN^{n,k,b}_{t}+\left(S_{t}+\delta^{n,k,a}_{t}\right)\,dN^{n,k,a}_{t}\right)$
$=-S_{t}\,dq_{t}+\sum_{\begin{subarray}{c}n\in\mathcal{N}\\ k\in\mathcal{K}\end{subarray}}\Delta^{k}\left(\delta^{n,k,b}_{t}\,dN^{n,k,b}_{t}+\delta^{n,k,a}_{t}\,dN^{n,k,a}_{t}\right),$
where $\delta^{n,k,b}_{t}:=S_{t}-S^{n,k,b}_{t}$ and $\delta^{n,k,a}_{t}:=S^{n,k,a}_{t}-S_{t}$ are the bid and ask quote offsets – that we refer to, by abuse of language, as the bid and ask quotes. They represent the markdown and markup relative to the reference price $S_{t}$.
2.2 The Optimization Problem
In the above discussion, we introduced the three processes central to most market making models: the reference price process $\left(S_{t}\right)_{t\geq 0}$, the inventory process $\left(q_{t}\right)_{t\geq 0}$, and the cash process $\left(X_{t}\right)_{t\geq 0}$.
Assuming that the market maker continuously marks their inventory to the reference price, the dynamics of the profit and loss process $\left(\mathrm{PnL}_{t}\right)_{t\geq 0}$ of the market maker can be computed from Eqs. (1), (2) and (4):
$\displaystyle d\mathrm{PnL}_{t}$ $\displaystyle=dX_{t}+S_{t}\,dq_{t}+q_{t-}\,dS_{t}+d\langle q_{t},S_{t}\rangle$
$\displaystyle=\sum_{\begin{subarray}{c}n\in\mathcal{N}\\ k\in\mathcal{K}\end{subarray}}\left(\big{(}\Delta^{k}\delta^{n,k,b}_{t}-(q_{t-}+\Delta^{k})\,\tilde{\zeta}^{n,k}(\delta^{n,k,b}_{t})\big{)}dN^{n,k,b}_{t}+\big{(}\Delta^{k}\delta^{n,k,a}_{t}+(q_{t-}-\Delta^{k})\,\tilde{\zeta}^{n,k}(\delta^{n,k,a}_{t})\big{)}dN^{n,k,a}_{t}\right)$
$\displaystyle\quad+q_{t-}\,\sigma\,dB_{t}+q_{t-}\sum_{n\in\mathcal{N}}\tilde{J}^{n}\left(\sum_{k\in\mathcal{K}}w^{n,k}\left(\delta^{n,k,a}_{t}-\delta^{n,k,b}_{t}\right)\right)dt.$
A classical optimization approach is then to consider a finite-horizon objective, where the market maker maximizes the expected PnL at a given time $T$, while penalizing inventory risk over the interval $[0,T]$. This leads to the following risk-adjusted expected PnL objective, common in the works of Cartea, Jaimungal, and coauthors (see e.g. [8]):
$\mathbb{E}\left[\mathrm{PnL}_{T}-\frac{1}{2}\gamma\int_{0}^{T}\sigma^{2}q_{t}^{2}\,dt\right],$
maximized over the family of predictable and bounded-from-below control processes $\left(\left(\delta_{t}^{n,k,b},\delta_{t}^{n,k,a}\right)_{t\geq 0}\right)_{(n,k)\in\mathcal{N}\times\mathcal{K}}$. See also [13] for a broader perspective on market making objectives.
In this paper, we instead adopt a stationary infinite-horizon perspective, [^6] In market making models, the focus naturally lies on the optimal behavior far from the terminal time (which is often arbitrary). In the analysis that follows, we use Taylor expansions to capture the first-order impact of informational risk, working directly in an infinite-horizon framework. Performing these expansions in a finite-horizon setting and then taking the limit as $T\to+\infty$ would lead to results with no financial relevance (this is the classical iterated limit problem). and consider the discounted objective:
$\mathbb{E}\left[\int_{0}^{+\infty}e^{-\rho t}\left(d\mathrm{PnL}_{t}-\frac{\gamma}{2}\sigma^{2}q_{t}^{2}\,dt\right)\right],$
where $\rho>0$ is a discount rate.
Using the form of intensities given by Eq. (3), this objective is equivalent to
$\mathbb{E}\left[\int_{0}^{+\infty}e^{-\rho t}\left(\sum_{n\in\mathcal{N}}\beta^{n}\left(q_{t-},\delta^{n,\cdot,b}_{t},\delta^{n,\cdot,a}_{t}\right)-\frac{\gamma}{2}\sigma^{2}q_{t}^{2}\right)dt\right],$ (5)
where for each $n\in\mathcal{N}$, the function $\beta^{n}$ is defined as:
$\displaystyle\beta^{n}\Big{(}q,\delta^{n,\cdot,b},\delta^{n,\cdot,a}\Big{)}$ $\displaystyle=q\,\tilde{J}^{n}\left(\sum_{k\in\mathcal{K}}w^{n,k}\left(\delta^{n,k,a}-\delta^{n,k,b}\right)\right)$
$\displaystyle\quad+\sum_{k\in\mathcal{K}}\Bigg{(}\Delta^{k}\,\Lambda^{n,k,b}\left(\delta^{n,k,b}\right)\left(\delta^{n,k,b}-\frac{q+\Delta^{k}}{\Delta^{k}}\tilde{\zeta}^{n,k}\left(\delta^{n,k,b}\right)\right)$
$\displaystyle\qquad\qquad+\Delta^{k}\,\Lambda^{n,k,a}\left(\delta^{n,k,a}\right)\left(\delta^{n,k,a}+\frac{q-\Delta^{k}}{\Delta^{k}}\tilde{\zeta}^{n,k}\left(\delta^{n,k,a}\right)\right)\Bigg{)}.$
3 Informational Risk: A First-Order Perspective
The aim of this section is to understand the influence of informational risk on the optimal behavior of a market maker. To this end, we analyze the optimal control problem associated with the objective function in Eq. (5). Rather than solving the problem fully, we aim to characterize the first-order effects of informational risk. We begin by presenting the Hamilton-Jacobi-Bellman (HJB) equation associated with the general model, then recall classical results in the absence of informational risk, and finally compute first-order adjustments.
3.1 The Hamilton-Jacobi-Bellman Equation
The HJB equation corresponding to the infinite horizon problem (5) is a nonlinear system for the value function $\vartheta:\Delta^{1}\mathbb{Z}\to\mathbb{R}$, given by:
$\displaystyle 0$ $\displaystyle=-\rho\,\vartheta(q)-\frac{1}{2}\gamma\sigma^{2}q^{2}+\sum_{n\in\mathcal{N}}\tilde{\mathcal{H}}^{n}\Bigl{[}\tilde{J}^{n},\tilde{\zeta}^{n,\cdot}\Bigr{]}\left(q,D^{\cdot}_{+}\vartheta(q),D^{\cdot}_{-}\vartheta(q)\right),$ (6)
where the finite-difference operators are defined, for all $k\in\mathcal{K}$, by
$D^{k}_{+}\vartheta(q):=\frac{\vartheta(q)-\vartheta(q+\Delta^{k})}{\Delta^{k}}\quad\text{and}\quad D^{k}_{-}\vartheta(q):=\frac{\vartheta(q)-\vartheta(q-\Delta^{k})}{\Delta^{k}},$
and for each $n\in\mathcal{N}$, the Hamiltonian $\tilde{\mathcal{H}}^{n}$ is given by
$\displaystyle\tilde{\mathcal{H}}^{n}\left[\tilde{J}^{n},\tilde{\zeta}^{n,\cdot}\right]\left(q,p^{\cdot,b},p^{\cdot,a}\right):=\ \sup_{\delta^{n,\cdot,b},\delta^{n,\cdot,a}}\Bigg{\{}\beta^{n}\left(q,\delta^{n,\cdot,b},\delta^{n,\cdot,a}\right)$
$\displaystyle\qquad\qquad\qquad\qquad\qquad-\sum_{k\in\mathcal{K}}\Delta^{k}\left(\Lambda^{n,k,b}(\delta^{n,k,b})\,p^{k,b}+\Lambda^{n,k,a}(\delta^{n,k,a})\,p^{k,a}\right)\Bigg{\}}.$
When they exist, we denote by $\tilde{\mathfrak{d}}^{n,\cdot,b/a*}(q,p^{\cdot,b},p^{\cdot,a})$ the maximizers in the above definition. The optimal quotes for tier $n$ in feedback form, denoted by $\mathfrak{d}^{n,\cdot,b/a*}(q)$, are then given by $\tilde{\mathfrak{d}}^{n,\cdot,b/a*}\left(q,D^{\cdot}_{+}\vartheta(q),D^{\cdot}_{-}\vartheta(q)\right)$.
Remark 3.
The value function $\vartheta$ and the optimal quotes depend on the discount rate $\rho$ whose value may appear arbitrary. When $\rho\to 0$, the value function blows up, but $\rho\vartheta$ converges towards a finite value called the ergodic constant of the problem. More importantly, finite differences of the value function, i.e. terms of the forms $D^{k}_{+}\vartheta(q)$ and $D^{k}_{-}\vartheta(q)$, converge, hence the convergence of optimal quotes when $\rho\to 0$.
3.2 The Baseline without Informational Risk
Solving Eq. (6) in full generality requires the introduction of assumptions on the functions $\tilde{J}^{\cdot}$ and $\tilde{\zeta}^{\cdot,\cdot}$. Since our objective is instead to investigate first-order informational effects, we begin with the classical baseline case where there is no informational risk – that is, we assume $\tilde{J}^{n}=0$ and $\tilde{\zeta}^{n,k}=0$ for all $(n,k)\in\mathcal{N}\times\mathcal{K}$. In this setting, the reference price evolves as a Brownian motion and the baseline value function $\theta:\Delta^{1}\mathbb{Z}\to\mathbb{R}$ solves the simpler HJB equation:
$0=\ -\rho\,\theta(q)-\frac{1}{2}\gamma\sigma^{2}q^{2}+\sum_{\begin{subarray}{c}n\in\mathcal{N}\\ k\in\mathcal{K}\end{subarray}}\Delta^{k}H^{n,k,b}\left(D^{k}_{+}\theta(q)\right)+\sum_{\begin{subarray}{c}n\in\mathcal{N}\\ k\in\mathcal{K}\end{subarray}}\Delta^{k}H^{n,k,a}\left(D^{k}_{-}\theta(q)\right),$
where the Hamiltonian functions are defined, for all $(n,k)\in\mathcal{N}\times\mathcal{K}$, by
$H^{n,k,b/a}(p):=\sup_{\delta\in\mathbb{R}}\Lambda^{n,k,b/a}(\delta)\,(\delta-p),$
and the associated maximizers [^7] They verify the simple ordinary differential equation $\tilde{\delta}^{n,k,b/a*^{\prime}}(p)=\frac{1}{c^{n,k,b/a}(\tilde{\delta}^{n,k,b/a*}(p))}$. are given by
$\tilde{\delta}^{n,k,b/a*}(p):=\text{argmax}_{\delta\in\mathbb{R}}\Lambda^{n,k,b/a}(\delta)\,(\delta-p)=\left(\Lambda^{n,k,b/a}\right)^{-1}\left(-{H^{n,k,b/a}}^{\prime}(p)\right).$
Given the baseline value function $\theta$, the baseline optimal quotes in feedback form are therefore the following, for all $(n,k)\in\mathcal{N}\times\mathcal{K}$:
$\delta^{n,k,b*}(q)=\tilde{\delta}^{n,k,b*}\left(D^{k}_{+}\theta(q)\right)\quad\text{and}\quad\delta^{n,k,a*}(q)=\tilde{\delta}^{n,k,a*}\left(D^{k}_{-}\theta(q)\right).$
3.3 First-Order Effects
3.3.1 Framework
In order to study the first-order effects of adverse selection and price reading, we introduce a perturbation parameter $\varepsilon>0$ and set
$\tilde{J}^{n}(x)=\varepsilon J^{n}(x)\quad\text{for all }n\in\mathcal{N}\qquad\text{and}\qquad\tilde{\zeta}^{n,k}(\delta)=\varepsilon\zeta^{n,k}(\delta)\quad\text{for all }(n,k)\in\mathcal{N}\times\mathcal{K}.$
For notational convenience, we define the $\varepsilon$-dependent Hamiltonian functions by
$\mathcal{H}^{n}\left(q,p^{\cdot,b},p^{\cdot,a},\varepsilon\right):=\ \tilde{\mathcal{H}}^{n}\left[\varepsilon J^{n},\varepsilon\zeta^{n,\cdot}\right]\left(q,p^{\cdot,b},p^{\cdot,a}\right)=\ \sup_{\delta^{n,\cdot,b},\delta^{n,\cdot,a}}\mathcal{B}^{n}\Big{(}q,p^{\cdot,b},p^{\cdot,a},\delta^{n,\cdot,b},\delta^{n,\cdot,a},\varepsilon\Big{)},$
where
$\displaystyle\mathcal{B}^{n}\Big{(}$ $\displaystyle q,p^{\cdot,b},p^{\cdot,a},\delta^{n,\cdot,b},\delta^{n,\cdot,a},\varepsilon\Big{)}:=\ q\,\varepsilon J^{n}\left(\sum_{k\in\mathcal{K}}w^{n,k}\left(\delta^{n,k,a}-\delta^{n,k,b}\right)\right)$
$\displaystyle+\sum_{k\in\mathcal{K}}\Delta^{k}\,\Lambda^{n,k,b}\left(\delta^{n,k,b}\right)\left(\delta^{n,k,b}-p^{k,b}-\frac{q+\Delta^{k}}{\Delta^{k}}\varepsilon\zeta^{n,k}\left(\delta^{n,k,b}\right)\right)$
$\displaystyle+\sum_{k\in\mathcal{K}}\Delta^{k}\,\Lambda^{n,k,a}\left(\delta^{n,k,a}\right)\left(\delta^{n,k,a}-p^{k,a}+\frac{q-\Delta^{k}}{\Delta^{k}}\varepsilon\zeta^{n,k}\left(\delta^{n,k,a}\right)\right).$
3.3.2 Value Function
We then seek a solution to Eq. (6) of the form:
$\vartheta(q)=\theta(q)+\varepsilon f(q)+o(\varepsilon),$
where $f$ captures the first-order impact of informational effects on the value function.
To compute the equation for $f$, we need the gradient of the Hamiltonian functions. Notice that:
$\nabla_{\left(\begin{smallmatrix}p^{\cdot,b}\\ p^{\cdot,a}\end{smallmatrix}\right)}\mathcal{H}^{n}\left(q,p^{\cdot,b},p^{\cdot,a},0\right)=-\left(\begin{matrix}\Delta^{\cdot}\,\Lambda^{n,\cdot,b}\left(\tilde{\delta}^{n,\cdot,b*}(p^{\cdot,b})\right)\\ \Delta^{\cdot}\,\Lambda^{n,\cdot,a}\left(\tilde{\delta}^{n,\cdot,a*}(p^{\cdot,a})\right)\end{matrix}\right)=\left(\begin{matrix}\Delta^{\cdot}\,H^{n,\cdot,b}\,{}^{\prime}\left(p^{\cdot,b}\right)\\ \Delta^{\cdot}\,H^{n,\cdot,a}\,{}^{\prime}\left(p^{\cdot,a}\right)\end{matrix}\right)$
and
$\displaystyle\partial_{\varepsilon}\mathcal{H}^{n}(q,p^{\cdot,b},p^{\cdot,a},0)$ $\displaystyle=qJ^{n}\left(\sum_{k\in\mathcal{K}}w^{n,k}\big{(}\tilde{\delta}^{n,k,a*}(p^{k,a})-\tilde{\delta}^{n,k,b*}(p^{k,b})\big{)}\right)-\sum_{k\in\mathcal{K}}(q+\Delta^{k})\left(\Lambda^{n,k,b}\zeta^{n,k}\right)\big{(}\tilde{\delta}^{n,k,b*}(p^{k,b})\big{)}$
$\displaystyle\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad+\sum_{k\in\mathcal{K}}(q-\Delta^{k})\left(\Lambda^{n,k,a}\zeta^{n,k}\right)\big{(}\tilde{\delta}^{n,k,a*}(p^{k,a})\big{)}.$
Injecting $\vartheta(q)=\theta(q)+\varepsilon f(q)+o(\varepsilon)$ into Eq. (6) and collecting $\mathcal{O}(\varepsilon)$ terms yields the linear equation:
$\displaystyle 0=$ $\displaystyle-\rho f(q)+\sum_{n\in\mathcal{N}}\Bigg{(}qJ^{n}\left(\sum_{k\in\mathcal{K}}w^{n,k}(\delta^{n,k,a*}(q)-\delta^{n,k,b*}(q))\right)$ (7)
$\displaystyle-\sum_{k\in\mathcal{K}}\Lambda^{n,k,b}(\delta^{n,k,b*}(q))\left(f(q)-f(q+\Delta^{k})+(q+\Delta^{k})\zeta^{n,k}(\delta^{n,k,b*}(q))\right)$
$\displaystyle-\sum_{k\in\mathcal{K}}\Lambda^{n,k,a}(\delta^{n,k,a*}(q))\left(f(q)-f(q-\Delta^{k})-(q-\Delta^{k})\zeta^{n,k}(\delta^{n,k,a*}(q))\right)\Bigg{)}.$
Using a Feynman-Kac representation, and defining $\delta^{n,k,b/a*}_{t}=\delta^{n,k,b/a*}(q_{t-})$, the solution can be written as a sum over tiers
$f(q)=\sum_{n\in\mathcal{N}}f^{n}(q),$
where, for each $n\in\mathcal{N}$,
$\displaystyle f^{n}(q)=\mathbb{E}\Bigg{[}\int_{0}^{\infty}e^{-\rho t}\Bigg{(}$ $\displaystyle q_{t-}J^{n}\Big{(}\sum_{k\in\mathcal{K}}w^{n,k}(\delta^{n,k,a*}_{t}-\delta^{n,k,b*}_{t})\Big{)}\,dt$
$\displaystyle-\sum_{k\in\mathcal{K}}\Big{(}(q_{t-}+\Delta^{k})\zeta^{n,k}(\delta^{n,k,b*}_{t})\,dN^{n,k,b}_{t}-(q_{t-}-\Delta^{k})\zeta^{n,k}(\delta^{n,k,a*}_{t})\,dN^{n,k,a}_{t}\Big{)}\Bigg{)}\,\bigg{|}\,q_{0}=q\Bigg{]}.$
In other words, $f$ measures the impact of informational risk on the risk-adjusted expected PnL if the market maker kept the quotes that were optimal in the absence of informational risk.
3.3.3 Optimal Quotes
The above first-order expansion quantifies the impact of informational risk on the risk-adjusted expected PnL of the market maker. In order to address the optimal behavior of the market maker, we also need a first-order expansion of the maximizers in the definition of the Hamiltonian functions.
Writing
$\tilde{\mathfrak{d}}^{n,k,b/a*}(q,p^{\cdot,b},p^{\cdot,a})=\tilde{\delta}^{n,k,b/a*}(p^{\cdot,b/a})+\varepsilon g^{n,k,b/a}(q,p^{\cdot,b},p^{\cdot,a})+o(\varepsilon),$
we have
$\displaystyle\nabla_{\left(\begin{smallmatrix}\delta^{n,\cdot,b}\\ \delta^{n,\cdot,a}\end{smallmatrix}\right)}\mathcal{B}^{n}\Big{(}$ $\displaystyle q,\,p^{\cdot,b},p^{\cdot,a},\tilde{\delta}^{n,\cdot,b*}(p^{\cdot,b})+\varepsilon g^{n,\cdot,b}(q,p^{\cdot,b},p^{\cdot,a}),\tilde{\delta}^{n,\cdot,a*}(p^{\cdot,a})+\varepsilon g^{n,\cdot,a}(q,p^{\cdot,b},p^{\cdot,a}),\varepsilon\Big{)}=o(\varepsilon).$
Expanding in $\varepsilon$, and using the first-order condition in the baseline case, we get
$\displaystyle\nabla^{2}_{\left(\begin{smallmatrix}\delta^{n,\cdot,b}\\ \delta^{n,\cdot,a}\end{smallmatrix}\right),\left(\begin{smallmatrix}\delta^{n,\cdot,b}\\ \delta^{n,\cdot,a}\end{smallmatrix}\right)}\mathcal{B}^{n}\Big{(}q,\,p^{\cdot,b},\,p^{\cdot,a},\,\tilde{\delta}^{n,\cdot,b*}(p^{\cdot,b}),\,\tilde{\delta}^{n,\cdot,a*}(p^{\cdot,b}),\,0\Big{)}\begin{pmatrix}g^{n,\cdot,b}(q,p^{\cdot,b},p^{\cdot,a})\\ g^{n,\cdot,a}(q,p^{\cdot,b},p^{\cdot,a})\end{pmatrix}$
[…]
A straightforward computation yields then
$g^{n,k,b}(q,p^{\cdot,b},p^{\cdot,a})=\frac{qw^{n,k}J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\tilde{\delta}^{n,j,a*}(p^{j,a})-\tilde{\delta}^{n,j,b*}(p^{j,b})\right)\right)+(q+\Delta^{k})\left(\Lambda^{n,k,b}\zeta^{n,k}\right)^{\prime}\!\left(\tilde{\delta}^{n,k,b*}(p^{k,b})\right)}{\Delta^{k}\Lambda^{n,k,b\,\prime}\!\left(\tilde{\delta}^{n,k,b*}(p^{k,b})\right)c^{n,k,b}(\tilde{\delta}^{n,k,b*}(p^{k,b}))}$
and
$g^{n,k,a}(q,p^{\cdot,b},p^{\cdot,a})=\frac{-qw^{n,k}J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\tilde{\delta}^{n,j,a*}(p^{j,a})-\tilde{\delta}^{n,j,b*}(p^{j,b})\right)\right)-(q-\Delta^{k})\left(\Lambda^{n,k,a}\zeta^{n,k}\right)^{\prime}\!\left(\tilde{\delta}^{n,k,a*}(p^{k,a})\right)}{\Delta^{k}\Lambda^{n,k,a\,\prime}\!\left(\tilde{\delta}^{n,k,a*}(p^{k,a})\right)c^{n,k,a}(\tilde{\delta}^{n,k,a*}(p^{k,a}))}.$
To get the optimal quotes in feedback form, we combine the above two first-order expansions to get
$\displaystyle\mathfrak{d}^{n,k,b*}(q)$ $\displaystyle=$ $\displaystyle\tilde{\mathfrak{d}}^{n,k,b*}(q,D^{\cdot}_{+}\vartheta(q),D^{\cdot}_{-}\vartheta(q))$
$\displaystyle=$ $\displaystyle\tilde{\delta}^{n,k,b*}(D^{k}_{+}\theta(q)+\varepsilon D^{k}_{+}f(q))+\varepsilon g^{n,k,b}(q,D^{\cdot}_{+}\theta(q)+\varepsilon D^{k}_{+}f(q),D^{\cdot}_{-}\theta(q)+\varepsilon D^{k}_{-}f(q))+o(\varepsilon)$
$\displaystyle=$ $\displaystyle\tilde{\delta}^{n,k,b*}(D^{k}_{+}\theta(q))+\varepsilon\tilde{\delta}^{n,k,b*^{\prime}}(D^{k}_{+}\theta(q))D^{k}_{+}f(q)+\varepsilon g^{n,k,b}(q,D^{\cdot}_{+}\theta(q),D^{\cdot}_{-}\theta(q))+o(\varepsilon)$
$\displaystyle=$ $\displaystyle\delta^{n,k,b*}(q)+\frac{\varepsilon}{c^{n,k,b}(\delta^{n,k,b*}(q))}\left(D^{k}_{+}f(q)+\frac{qw^{n,k}J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\delta^{n,j,a*}(q)-\delta^{n,j,b*}(q)\right)\right)}{{\Delta^{k}}\Lambda^{n,k,b\,\prime}\!\left(\delta^{n,k,b*}(q)\right)}\right$.
$\displaystyle\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.+\frac{(q+\Delta^{k})\left(\Lambda^{n,k,b}\zeta^{n,k}\right)^{\prime}\!\left(\delta^{n,k,b*}(q)\right)}{{\Delta^{k}}\Lambda^{n,k,b\,\prime}\!\left(\delta^{n,k,b*}(q)\right)}\right)+o(\varepsilon)$
and similarly
$\displaystyle\mathfrak{d}^{n,k,a*}(q)$ $\displaystyle=$ $\displaystyle\delta^{n,k,a*}(q)+\frac{\varepsilon}{c^{n,k,a}(\delta^{n,k,a*}(q))}\left(D^{k}_{-}f(q)-\frac{qw^{n,k}J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\delta^{n,j,a*}(q)-\delta^{n,j,b*}(q)\right)\right)}{{\Delta^{k}}\Lambda^{n,k,a\,\prime}\!\left(\delta^{n,k,a*}(q)\right)}\right$.
$\displaystyle\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.-\frac{(q-\Delta^{k})\left(\Lambda^{n,k,a}\zeta^{n,k}\right)^{\prime}\!\left(\delta^{n,k,a*}(q)\right)}{{\Delta^{k}}\Lambda^{n,k,a\,\prime}\!\left(\delta^{n,k,a*}(q)\right)}\right)+o(\varepsilon)$
These expressions highlight that the influence of informational risk on optimal quotes can be decomposed into two components. The first component arises from the terms $D^{k}_{\pm}f$, which are finite differences of the function $f$. By definition, these terms reflect the marginal impact of informational risk on the risk-adjusted expected PnL, assuming the market maker does not update their quotes. As such, they aggregate the effects of informational risk across all tiers and are identical for all $n$. The second component is represented by the terms $g^{n,k,b/a}$, which quantify the impact of informational risk on the greedy quotes derived from the baseline value function in the absence of informational risk. Unlike the $D^{k}_{\pm}f$ terms, the $g^{n,k,b/a}$ terms are specific to the considered tier and trade size.
These two components may push optimal quotes in opposite directions, and their net effect typically requires numerical evaluation. Fortunately, this computation is straightforward, as it relies only on the optimal quotes in the baseline setting and the function $f$, which solves the infinite linear system (7) that can meaningfully be truncated. To shed further light on the mechanisms at play, we now turn to a class of closed-form approximations.
4 Quadratic Hamiltonian Approximation
In the algorithmic market making literature, closed-form approximations of the value function and optimal quotes are frequently used. The standard approach (see [6]) consists in replacing the HJB equation associated with the control problem by an alternative HJB equation in which the Hamiltonian functions are approximated to second order. The resulting modified HJB equation admits a solution that is quadratic in the inventory variable, with coefficients computable through a Riccati equation and a linear equation. [^8] While we focus on the single-asset case in this article, it is worth emphasizing that this methodology extends naturally to multi-asset market making, thereby offering an effective remedy to the curse of dimensionality.
4.1 Towards Quadratic Hamiltonian Functions
In this section, we expand the Hamiltonian functions to second order with respect to $\left(p^{\cdot,b},p^{\cdot,a},\varepsilon\right)$ around the point $(0,0,0)$.
From Section 3.3, we have the following first-order derivatives:
$\displaystyle\nabla_{\left(\begin{smallmatrix}p^{\cdot,b}\\ p^{\cdot,a}\end{smallmatrix}\right)}\mathcal{H}^{n}\left(q,0,0,0\right)$ $\displaystyle=\left(\begin{matrix}\Delta^{\cdot}\,H^{n,\cdot,b}\,{}^{\prime}(0)\\ \Delta^{\cdot}\,H^{n,\cdot,a}\,{}^{\prime}(0)\end{matrix}\right),$
and
$\displaystyle\partial_{\varepsilon}\mathcal{H}^{n}(q,0,0,0)$ $\displaystyle=q\,J^{n}\left(\sum_{k\in\mathcal{K}}w^{n,k}\left(\tilde{\delta}^{n,k,a*}(0)-\tilde{\delta}^{n,k,b*}(0)\right)\right)-\sum_{k\in\mathcal{K}}(q+\Delta^{k})\,\left(\Lambda^{n,k,b}\zeta^{n,k}\right)\left(\tilde{\delta}^{n,k,b*}(0)\right)$
$\displaystyle\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad+\sum_{k\in\mathcal{K}}(q-\Delta^{k})\,\left(\Lambda^{n,k,a}\zeta^{n,k}\right)\left(\tilde{\delta}^{n,k,a*}(0)\right)\,.$
Remark 4.
For all $(n,k)\in\mathcal{N}\times\mathcal{K}$, the quotes $\tilde{\delta}^{n,k,b*}(0)$ and $\tilde{\delta}^{n,k,a*}(0)$ correspond to the myopic quotes that a market maker with no risk aversion would set in the baseline case without informational risk.
For the second-order derivatives, we have
$\displaystyle\nabla^{2}_{\left(\begin{smallmatrix}p^{\cdot,b}\\ p^{\cdot,a}\\ \varepsilon\end{smallmatrix}\right),\left(\begin{smallmatrix}p^{\cdot,b}\\ p^{\cdot,a}\\ \varepsilon\end{smallmatrix}\right)}\mathcal{H}^{n}\left(q,0,0,0\right)$
$\displaystyle=$ $\displaystyle\nabla^{2}_{\left(\begin{smallmatrix}p^{\cdot,b}\\ p^{\cdot,a}\\ \varepsilon\end{smallmatrix}\right),\left(\begin{smallmatrix}p^{\cdot,b}\\ p^{\cdot,a}\\ \varepsilon\end{smallmatrix}\right)}\mathcal{B}^{n}(\bar{x}^{n})-\nabla^{2}_{\left(\begin{smallmatrix}p^{\cdot,b}\\ p^{\cdot,a}\\ \varepsilon\end{smallmatrix}\right),\left(\begin{smallmatrix}\delta^{\cdot,b}\\ \delta^{\cdot,a}\end{smallmatrix}\right)}\mathcal{B}^{n}(\bar{x}^{n})\left(\nabla^{2}_{\left(\begin{smallmatrix}\delta^{\cdot,b}\\ \delta^{\cdot,a}\end{smallmatrix}\right),\left(\begin{smallmatrix}\delta^{\cdot,b}\\ \delta^{\cdot,a}\end{smallmatrix}\right)}\mathcal{B}^{n}(\bar{x}^{n})\right)^{-1}\nabla^{2}_{\left(\begin{smallmatrix}\delta^{\cdot,b}\\ \delta^{\cdot,a}\end{smallmatrix}\right),\left(\begin{smallmatrix}p^{\cdot,b}\\ p^{\cdot,a}\\ \varepsilon\end{smallmatrix}\right)}\mathcal{B}^{n}(\bar{x}^{n})$
where
$\bar{x}^{n}:=\left(q,\,0,\,0,\,\tilde{\delta}^{n,\cdot,b*}(0),\,\tilde{\delta}^{n,\cdot,a*}(0),\,0\right).$
Now, we easily see that
$\nabla^{2}_{\left(\begin{smallmatrix}p^{\cdot,b}\\ p^{\cdot,a}\\ \varepsilon\end{smallmatrix}\right),\left(\begin{smallmatrix}p^{\cdot,b}\\ p^{\cdot,a}\\ \varepsilon\end{smallmatrix}\right)}\mathcal{B}^{n}\left(\bar{x}^{n}\right)=0,$
and that
$\nabla^{2}_{\left(\begin{smallmatrix}\delta^{\cdot,b}\\ \delta^{\cdot,a}\end{smallmatrix}\right),\left(\begin{smallmatrix}\delta^{\cdot,b}\\ \delta^{\cdot,a}\end{smallmatrix}\right)}\mathcal{B}^{n}\left(\bar{x}^{n}\right)=\operatorname{diag}\left(-\Delta^{\cdot}\dfrac{\Lambda^{n,\cdot,b\,\prime}(\tilde{\delta}^{n,\cdot,b*}(0))^{2}}{H^{n,\cdot,b\,\prime\prime}(0)},-\Delta^{\cdot}\dfrac{\Lambda^{n,\cdot,a\,\prime}(\tilde{\delta}^{n,\cdot,a*}(0))^{2}}{H^{n,\cdot,a\,\prime\prime}(0)}\right)$
while
$\displaystyle\nabla^{2}_{\left(\begin{smallmatrix}p^{\cdot,b}\\ p^{\cdot,a}\\ \varepsilon\end{smallmatrix}\right),\left(\begin{smallmatrix}\delta^{\cdot,b}\\ \delta^{\cdot,a}\end{smallmatrix}\right)}\mathcal{B}^{n}\left(\bar{x}^{n}\right)$ $\displaystyle=\begin{pmatrix}\operatorname{diag}\left(-\Delta^{\cdot}\,\Lambda^{n,\cdot,b\,\prime}\!\left(\tilde{\delta}^{n,\cdot,b*}(0)\right)\right)&0\\[4.30554pt] 0&\operatorname{diag}\left(-\Delta^{\cdot}\,\Lambda^{n,\cdot,a\,\prime}\!\left(\tilde{\delta}^{n,\cdot,a*}(0)\right)\right)\\[4.30554pt] \eta^{\cdot\top}&\iota^{\cdot\top}\end{pmatrix}$,
where, for all $k\in\mathcal{K}$,
$\eta^{k}:=-q\,w^{n,k}\,J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\tilde{\delta}^{n,j,a*}(0)-\tilde{\delta}^{n,j,b*}(0)\right)\right)-(q+\Delta^{k})\,\left(\Lambda^{n,k,b}\zeta^{n,k}\right)^{\prime}\!\left(\tilde{\delta}^{n,k,b*}(0)\right),$
$\iota^{k}:=\phantom{-}q\,w^{n,k}\,J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\tilde{\delta}^{n,j,a*}(0)-\tilde{\delta}^{n,j,b*}(0)\right)\right)+(q-\Delta^{k})\,\left(\Lambda^{n,k,a}\zeta^{n,k}\right)^{\prime}\!\left(\tilde{\delta}^{n,k,a*}(0)\right).$
Therefore,
$\nabla^{2}_{\left(\begin{smallmatrix}p^{\cdot,b}\\ p^{\cdot,a}\\ \varepsilon\end{smallmatrix}\right),\left(\begin{smallmatrix}p^{\cdot,b}\\ p^{\cdot,a}\\ \varepsilon\end{smallmatrix}\right)}\mathcal{H}^{n}\left(q,0,0,0\right)$
$\begin{pmatrix}\Delta^{\cdot}\,H^{n,\cdot,b\,\prime\prime}(0)&0&-\dfrac{H^{n,\cdot,b\,\prime\prime}(0)}{\Lambda^{n,\cdot,b\,\prime}\left(\tilde{\delta}^{n,\cdot,b*}(0)\right)}\,\eta^{\cdot}\\[6.45831pt] 0&\Delta^{\cdot}\,H^{n,\cdot,a\,\prime\prime}(0)&-\dfrac{H^{n,\cdot,a\,\prime\prime}(0)}{\Lambda^{n,\cdot,a\,\prime}\left(\tilde{\delta}^{n,\cdot,a*}(0)\right)}\,\iota^{\cdot}\\[6.45831pt] -\dfrac{H^{n,\cdot,b\,\prime\prime}(0)}{\Lambda^{n,\cdot,b\,\prime}\left(\tilde{\delta}^{n,\cdot,b*}(0)\right)}\,\eta^{\cdot\top}&-\dfrac{H^{n,\cdot,a\,\prime\prime}(0)}{\Lambda^{n,\cdot,a\,\prime}\left(\tilde{\delta}^{n,\cdot,a*}(0)\right)}\,\iota^{\cdot\top}&\begin{aligned} -\sum_{k\in\mathcal{K}}&\left[\dfrac{H^{n,k,b\,\prime\prime}(0)}{\Delta^{k}\,\Lambda^{n,k,b\,\prime}\left(\tilde{\delta}^{n,k,b*}(0)\right)^{2}}\,{\eta^{k}}^{2}\right.\\ &\left.+\dfrac{H^{n,k,a\,\prime\prime}(0)}{\Delta^{k}\,\Lambda^{n,k,a\,\prime}\left(\tilde{\delta}^{n,k,a*}(0)\right)^{2}}\,{\iota^{k}}^{2}\right]\end{aligned}\end{pmatrix}$.
4.2 HJB Equation and Value Function
We now consider, instead of Eq. (6), the following equation:
$\displaystyle 0=\$$\displaystyle-\rho\,\widehat{\vartheta}(q)-\frac{1}{2}\gamma\sigma^{2}q^{2}+\sum_{n\in\mathcal{N}}\mathcal{H}^{n}\left(q,0,0,0\right)+\sum_{n\in\mathcal{N}}\nabla_{\left(\begin{smallmatrix}p^{\cdot,b}\\ p^{\cdot,a}\\ \varepsilon\end{smallmatrix}\right)}\mathcal{H}^{n}\left(q,0,0,0\right)^{\top}\begin{pmatrix}D^{\cdot}_{+}\widehat{\vartheta}(q)\\ D^{\cdot}_{-}\widehat{\vartheta}(q)\\ \varepsilon\end{pmatrix}$
$\displaystyle+\frac{1}{2}\sum_{n\in\mathcal{N}}\begin{pmatrix}D^{\cdot}_{+}\widehat{\vartheta}(q)\\ D^{\cdot}_{-}\widehat{\vartheta}(q)\\ \varepsilon\end{pmatrix}^{\top}\nabla^{2}_{\left(\begin{smallmatrix}p^{\cdot,b}\\ p^{\cdot,a}\\ \varepsilon\end{smallmatrix}\right),\left(\begin{smallmatrix}p^{\cdot,b}\\ p^{\cdot,a}\\ \varepsilon\end{smallmatrix}\right)}\mathcal{H}^{n}\left(q,0,0,0\right)\begin{pmatrix}D^{\cdot}_{+}\widehat{\vartheta}(q)\\ D^{\cdot}_{-}\widehat{\vartheta}(q)\\ \varepsilon\end{pmatrix}$. (8)
A straightforward computation shows that the solution $\widehat{\vartheta}$ to Eq. (4.2) is of the form $\widehat{\vartheta}(q)=-A_{\varepsilon}q^{2}-B_{\varepsilon}q-C_{\varepsilon}$ with $A_{\varepsilon}>0$ and $B_{\varepsilon}$ satisfying [^9] We do no report any equation for $C_{\varepsilon}$ as it does not play any role for the optimal quotes. Note that, when $\rho\to 0$, $C_{\varepsilon}$ typically blows up, but $\rho C_{\varepsilon}$ converges towards a finite value. the following equations:
$\rho A_{\varepsilon}-\frac{1}{2}\gamma\sigma^{2}+2A_{\varepsilon}^{2}F^{2,1}_{+}+2\varepsilon A_{\varepsilon}\Sigma_{+}^{1,0}=O(\varepsilon^{2}),$ (9)
$\rho B_{\varepsilon}+2A_{\varepsilon}B_{\varepsilon}F^{2,1}_{+}+2A_{\varepsilon}^{2}F^{2,2}_{-}+2A_{\varepsilon}F^{1,1}_{-}+\varepsilon B_{\varepsilon}\Sigma^{1,0}_{+}+\varepsilon A_{\varepsilon}\Sigma^{3,1}_{-}+\varepsilon S=\mathcal{O}(\varepsilon^{2}),$ (10)
where
$F^{i,l}_{\pm}=\sum_{\begin{subarray}{c}n\in\mathcal{N}\\ k\in\mathcal{K}\end{subarray}}\left(\Delta^{k}\right)^{l}\left({H^{n,k,b}}^{(i)}(0)\pm{H^{n,k,a}}^{(i)}(0)\right),$
$\displaystyle\Sigma^{r,l}_{\pm}$ $\displaystyle=\sum_{\begin{subarray}{c}n\in\mathcal{N}\\ k\in\mathcal{K}\end{subarray}}\left(\Delta^{k}\right)^{l}\left[\frac{H^{n,k,b\,\prime\prime}(0)}{\Lambda^{n,k,b\,\prime}\left(\tilde{\delta}^{n,k,b*}(0)\right)}\left(w^{n,k}J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}(\tilde{\delta}^{n,j,a*}(0)-\tilde{\delta}^{n,j,b*}(0))\right)+r\left(\Lambda^{n,k,b}\zeta^{n,k}\right)^{\prime}\!\left(\tilde{\delta}^{n,k,b*}(0)\right)\right)\right$.
$\displaystyle\left.\quad\pm\frac{H^{n,k,a\,\prime\prime}(0)}{\Lambda^{n,k,a\,\prime}\left(\tilde{\delta}^{n,k,a*}(0)\right)}\left(w^{n,k}J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}(\tilde{\delta}^{n,j,a*}(0)-\tilde{\delta}^{n,j,b*}(0))\right)+r\left(\Lambda^{n,k,a}\zeta^{n,k}\right)^{\prime}\!\left(\tilde{\delta}^{n,k,a*}(0)\right)\right)\right],$
and
$S=\sum_{n\in\mathcal{N}}\left[J^{n}\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\tilde{\delta}^{n,j,a*}(0)-\tilde{\delta}^{n,j,b*}(0)\right)\right)-\sum_{k\in\mathcal{K}}\left(\left(\Lambda^{n,k,b}\zeta^{n,k}\right)\!\left(\tilde{\delta}^{n,k,b*}(0)\right)-\left(\Lambda^{n,k,a}\zeta^{n,k}\right)\!\left(\tilde{\delta}^{n,k,a*}(0)\right)\right)\right].$
From Eqs. (9) and (10), we get
$A_{\varepsilon}=A_{0}+\varepsilon A_{0}^{\prime}+o(\varepsilon)\qquad\text{and}\qquad B_{\varepsilon}=B_{0}+\varepsilon B_{0}^{\prime}+o(\varepsilon),$
where
$A_{0}=\frac{-\rho+\sqrt{\rho^{2}+4\gamma\sigma^{2}F_{+}^{2,1}}}{4F_{+}^{2,1}}\qquad\text{and}\qquad B_{0}=-\frac{2A_{0}}{\rho+2A_{0}F_{+}^{2,1}}\left(A_{0}F_{-}^{2,2}+D_{-}^{1,1}\right),$
and
$A_{0}^{\prime}=-\frac{2A_{0}\Sigma_{+}^{1,0}}{\rho+4A_{0}F_{+}^{2,1}}\text{ and }B_{0}^{\prime}=-\frac{1}{\rho+2A_{0}F_{+}^{2,1}}\left(B_{0}\Sigma^{1,0}_{+}+A_{0}\Sigma^{3,1}_{-}+S-\frac{4A_{0}\Sigma_{+}^{1,0}}{\rho+4A_{0}F_{+}^{2,1}}\left(B_{0}F^{2,1}_{+}+2A_{0}F^{2,2}_{-}+F^{1,1}_{-}\right)\right).$
Therefore, if $\widehat{\vartheta}_{0}(q)=-A_{0}q^{2}-B_{0}q$ and $\widehat{f}(q)=-A_{0}^{\prime}q^{2}-B_{0}^{\prime}q$, then our approximation of the value function using quadratic approximations of Hamiltonian functions is (up to an additive constant) given by:
$\widehat{\vartheta}(q)=\widehat{\vartheta}_{0}(q)+\varepsilon\widehat{f}(q)+o(\varepsilon).$
4.3 Optimal Quotes
If the greedy quotes in feedback form associated with $\widehat{\vartheta}_{0}$, i.e. in the case of no informational risk under the quadratic approximation, are denoted by
$\widehat{\delta}^{n,k,b*}(q):=\tilde{\delta}^{n,k,b*}\left(D^{k}_{+}\widehat{\vartheta}_{0}(q)\right)\quad\text{and}\quad\widehat{\delta}^{n,k,a*}(q):=\tilde{\delta}^{n,k,a*}\left(D^{k}_{-}\widehat{\vartheta}_{0}(q)\right),$
then, as in Section 3.3.3, the first-order effects of informational risk on optimal quotes in feedback form are given by
$\displaystyle\widehat{\mathfrak{d}}^{n,k,b*}(q)$ $\displaystyle=$ $\displaystyle\widehat{\delta}^{n,k,b*}(q)+\frac{\varepsilon}{c^{n,k,b}\left(\widehat{\delta}^{n,k,b*}(q)\right)}\left(D^{k}_{+}\widehat{f}(q)+\frac{qw^{n,k}J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\widehat{\delta}^{n,j,a*}(q)-\widehat{\delta}^{n,j,b*}(q)\right)\right)}{{\Delta^{k}}\Lambda^{n,k,b\,\prime}\!\left(\widehat{\delta}^{n,k,b*}(q)\right)}\right$.
$\displaystyle\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.+\frac{(q+\Delta^{k})\left(\Lambda^{n,k,b}\zeta^{n,k}\right)^{\prime}\!\left(\widehat{\delta}^{n,k,b*}(q)\right)}{{\Delta^{k}}\Lambda^{n,k,b\,\prime}\!\left(\widehat{\delta}^{n,k,b*}(q)\right)}\right)+o(\varepsilon)$
and
$\displaystyle\widehat{\mathfrak{d}}^{n,k,a*}(q)$ $\displaystyle=$ $\displaystyle\widehat{\delta}^{n,k,a*}(q)+\frac{\varepsilon}{c^{n,k,a}\left(\widehat{\delta}^{n,k,a*}(q)\right)}\left(D^{k}_{-}\widehat{f}(q)-\frac{qw^{n,k}J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\widehat{\delta}^{n,j,a*}(q)-\widehat{\delta}^{n,j,b*}(q)\right)\right)}{{\Delta^{k}}\Lambda^{n,k,a\,\prime}\!\left(\widehat{\delta}^{n,k,a*}(q)\right)}\right$.
$\displaystyle\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.-\frac{(q-\Delta^{k})\left(\Lambda^{n,k,a}\zeta^{n,k}\right)^{\prime}\!\left(\widehat{\delta}^{n,k,a*}(q)\right)}{{\Delta^{k}}\Lambda^{n,k,a\,\prime}\!\left(\widehat{\delta}^{n,k,a*}(q)\right)}\right)+o(\varepsilon).$
4.4 Analysis in the Symmetric Case
Given that $\widehat{f}$ is available in closed form, the above expressions are easier to handle than those in Section 3.3.3, which allows us to examine the underlying mechanisms in detail. For simplicity, we focus on a symmetric setting with identical bid- and ask-side functions, in the limiting case where $\rho\to 0$.
In this case indeed, we have $B_{\varepsilon}=B^{\prime}_{\varepsilon}=0$ and the first-order adjustment to the value function takes the form $\widehat{f}(q)=\frac{1}{2}\frac{\Sigma_{+}^{1,0}}{F_{+}^{2,1}}\,q^{2}$ (up to an additive constant) where [^10] We removed the $b$ and $a$ superscripts because of our assumption of symmetry.
$\Sigma^{1,0}_{+}=2\sum_{\begin{subarray}{c}m\in\mathcal{N}\\ j\in\mathcal{K}\end{subarray}}\frac{H^{m,j\,\prime\prime}(0)}{\Lambda^{m,j\,\prime}\left(\tilde{\delta}^{m,j*}(0)\right)}\left(w^{m,j}J^{m\,\prime}\!\left(0\right)+\left(\Lambda^{m,j}\zeta^{m,j}\right)^{\prime}\left(\tilde{\delta}^{m,j*}(0)\right)\right)\quad\text{and}\quad F^{2,1}_{+}=2\sum_{\begin{subarray}{c}m\in\mathcal{N}\\ j\in\mathcal{K}\end{subarray}}\Delta^{j}H^{m,j\,\prime\prime}(0).$
Because $J^{n}$ is non-decreasing, price reading contributes to increasing the concavity of the value function. The intuition is as follows: in the absence of informational risk, the market maker skews quotes to attract offsetting flows and deter trades that would increase the absolute value of the inventory. Skew sniffers detect this behavior and push the price in a direction unfavorable to the market maker, thereby amplifying inventory risk.
In contrast, the impact of adverse selection on the value function requires a more nuanced discussion. Each tier $n\in\mathcal{N}$ and size index $k\in\mathcal{K}$ contributes either positively or negatively to the change in concavity of the value function, depending on the sign of $\left(\zeta^{n,k}\Lambda^{n,k}\right)^{\prime}$. The sign of this term is determined by the nature of adverse selection within each tier.
Let us first consider a tier where clients trade on strong signals and would execute in the direction of their signal at virtually any reasonable price given prevailing market conditions. Such behavior is plausible for clients acting on mid- or long-term signals. In this case, the slopes of the corresponding $\zeta$ functions are insufficiently steep to compensate the decrease of intensity functions, resulting in a reduction in the concavity of the value function. The mechanism is as follows: if the market maker holds a positive (resp. negative) inventory, then by skewing – through a decrease (resp. increase) in both bid and ask prices – they raise the absolute and relative probability of executing a sell (resp. buy) order that moves the inventory closer to zero and therefore the likelihood that adverse selection – whose dependence on quotes is mild – induces a price increase (resp. decrease), thereby artificially increasing the value of the remaining inventory. Although adverse selection reduces the constant term in the approximation of the value function, this mechanism tends to flatten it, thus lowering its concavity.
Conversely, consider a tier where adverse selection predominantly arises when clients execute trades at prices set by the market maker with large offsets – typically the case of clients acting on short-term signals with a technological speed advantage. In such a scenario, the corresponding $\zeta$ functions may exhibit steep slopes, and the products $\zeta^{n,k}\Lambda^{n,k}$ may cease to be decreasing. This contributes to an increase in the concavity of the value function.
This analysis of the concavity of the value function provides valuable insight into the first component – common to all tiers – describing the impact of informational risk on optimal quotes. This component is proportional (with a positive sign on the bid side and a negative sign on the ask side) to the terms
$D^{k}_{\pm}\widehat{f}(q)=-\frac{\Sigma_{+}^{1,0}}{F_{+}^{2,1}}\left(\pm q+\frac{\Delta^{k}}{2}\right).$
When the inventory is zero, the preceding discussion justifies that this component leads to spread widening in the presence of price reading: the market maker compensates for the added informational risk by increasing the profit margin per trade. When the inventory is nonzero, the effect of price reading on the value function encourages the market maker to skew more aggressively, in order to shorten the time spent with a long or short position. The case of adverse selection is more nuanced, as it depends on the cumulative effect of the different tiers and sizes.
The second component of the optimal quotes, which is tier-specific and relates to the local adjustments of greedy quotes, consists of two terms: one related to price reading and the other to adverse selection.
Focusing first on price reading, the corresponding term in the quote offset has the sign of $-q$ on the bid side and $+q$ on the ask side. In particular, it vanishes when the inventory is zero and contributes to reducing the existing skew when the inventory is nonzero. This effect reflects the market maker’s incentive to reveal less information through their quotes in order to mitigate the impact of skew sniffers.
Taken together, the two components of the impact of price reading on optimal quotes lead to a subtle trade-off. When the inventory is zero, the second component vanishes and only the first component remains: a widening of the bid-ask spread to account for informational risk. When the inventory is nonzero, the first component encourages stronger skewing to reduce the time spent in unbalanced positions, while the second component pushes in the opposite direction, urging the market maker to reduce the skew in order to avoid revealing inventory information to skew sniffers. The second effect can dominate, for instance, in tiers with a large proportion of skew sniffers, especially when these clients trade infrequently but are highly sensitive to revealed information.
The case of adverse selection once again depends on the relative shape of the $\zeta$ and intensity functions. When adverse selection depends only mildly on the quotes, the second component affects the quote offsets according to the sign of $q+\Delta^{k}$ for bid transactions of size $\Delta^{k}$ and $-q+\Delta^{k}$ for ask transactions of the same size. In such cases, the $\Delta^{k}$ terms contribute to widening the spread, while the $q$ terms encourage additional skewing – opposite to the incentives from the first component. The reasoning is reversed when adverse selection primarily occurs at seemingly unfavorable prices, that is, when informed clients exploit short-term mispricings. However, the second component still pushes in the opposite direction to the first. In both scenarios, the second component may dominate the first in tiers with highly informed clients, especially when trades have a strong impact on prices. However, unlike in the case of price reading, liquidity – through the amplitude of $\Lambda^{n,k}$ – does not play a clear role in this mechanism.
5 Numerical Examples
5.1 Functional Forms
We now provide illustrations in the specific case
$\displaystyle\Lambda^{n,k}(\delta):=\Lambda^{n,k,b}(\delta)=\Lambda^{n,k,a}(\delta)=\Lambda_{0}^{n,k}e^{-\kappa^{n,k}\delta},\qquad J^{n}(x)=x\quad\text{and}\quad\zeta^{n,k}(\delta)=\alpha^{n,k}e^{\beta^{n,k}\delta},$ (11)
where for all $n\in\mathcal{N}$ and $k\in\mathcal{K}$, the parameters $\Lambda_{0}^{n,k},\kappa^{n,k},\alpha^{n,k},\beta^{n,k}$ are positive.
For such exponential intensities, we have
$c^{n,k,b/a*}=1,\qquad\tilde{\delta}^{n,k*}(p)=p+\frac{1}{\kappa^{n,k}}\quad\text{and}\quad H^{n,k}(p)=\frac{\Lambda_{0}^{n,k}}{\kappa^{n,k}}e^{-1-\kappa^{n,k}p},$
so that, in particular, $H^{n,k\,\prime\prime}(0)=\Lambda_{0}^{n,k}\kappa^{n,k}e^{-1}$.
Baseline pricing without informational risk in the quadratic approximation case is given by
$\widehat{\delta}^{n,k,b/a*}(q)=\frac{1}{\kappa^{n,k}}+\sigma\sqrt{\frac{\gamma e}{2\sum_{\begin{subarray}{c}m\in\mathcal{N}\\ j\in\mathcal{K}\end{subarray}}\Delta^{j}\Lambda_{0}^{m,j}\kappa^{m,j}}}\left(\pm q+\frac{\Delta^{k}}{2}\right).$
5.2 Price Reading
Let us consider the case of price reading only. [^11] We consider price reading and adverse selection separately in order to disentangle their respective effects. This separation is without loss of generality in our framework, which is restricted to first-order effects. We have
$D^{k}_{\pm}\widehat{f}(q)=\frac{e\sum_{\begin{subarray}{c}m\in\mathcal{N}\\ j\in\mathcal{K}\end{subarray}}w^{m,j}}{\sum_{\begin{subarray}{c}m\in\mathcal{N}\\ j\in\mathcal{K}\end{subarray}}\Delta^{j}\Lambda_{0}^{m,j}\kappa^{m,j}}\left(\pm q+\frac{\Delta^{k}}{2}\right),$
and therefore the quote adjustment is given by
$\frac{\widehat{\mathfrak{d}}^{n,k,b/a*}(q)-\widehat{\delta}^{n,k,b/a*}(q)}{\varepsilon}=\frac{e\sum_{\begin{subarray}{c}m\in\mathcal{N}\\ j\in\mathcal{K}\end{subarray}}w^{m,j}}{\sum_{\begin{subarray}{c}m\in\mathcal{N}\\ j\in\mathcal{K}\end{subarray}}\Delta^{j}\Lambda_{0}^{m,j}\kappa^{m,j}}\left(\pm q+\frac{\Delta^{k}}{2}\right)\mp\frac{qw^{n,k}}{\Delta^{k}\kappa^{n,k}}\frac{1}{\Lambda^{n,k}\left(\widehat{\delta}^{n,k,b/a*}(q)\right)}$
As discussed earlier, the first term captures the global response to the price reading risk, for trades of size $\Delta^{k}$, uniformly across all tiers, while the second term is tier-specific. This second term is non-linear in inventory and is highly sensitive to the presence and nature of skew sniffers in the tier: a market maker can tolerate skew sniffers with high trading activity because they contribute to global inventory risk reduction, but it is meaningless to show prices to skew sniffers who do not trade (indeed, the term diverges as $\Lambda_{0}^{n,k}\to 0$).
As an illustration, consider a standard size ladder of $(1,2,5,10,20,50)$ million notional and two tiers of equal baseline intensity, with $\Lambda_{0}^{n,k}=(1000,400,300,200,60,40)$ (in day-1 units) and $\kappa^{n,k}=3$ bp-1. This corresponds to a relatively liquid currency pair with a daily turnover exceeding 6 billion notional and a top-of-book spread of approximately $0.7$ bp, assuming a daily volatility of $100$ bp. In the first tier, we assume no price reading, i.e. $w^{1,k}=0$, whereas in the second tier we assume the presence of skew sniffers with $w^{2,k}=1$.
Figure 1: Optimal bid quote adjustment, $\varepsilon^{-1}\left(\widehat{\mathfrak{d}}^{n,k,b*}(q)-\widehat{\delta}^{n,k,b*}(q)\right)$ for the two tiers (we illustrate the first and the fourth sizes). Risk aversion $\gamma=10^{-5}$ bp-1 M-1. Other parameters are defined in the text.
Figure 2: Optimal bid quote adjustment, $\varepsilon^{-1}\left(\widehat{\mathfrak{d}}^{n,k,b*}(q)-\widehat{\delta}^{n,k,b*}(q)\right)$ for the two tiers (we illustrate the first and the fourth sizes). Risk aversion $\gamma=10^{-5}$ bp-1 M-1. $\Lambda_{0}^{1,k}=(1500,600,450,300,90,60)$, $\Lambda_{0}^{2,k}=(500,200,150,100,30,20)$. Other parameters are defined in the text.
As shown in Figure 1, [^12] We only plot bid quote adjustments because ask quote adjustments are symmetric. optimal pricing for the first tier reflects the global increase in informational risk, i.e. inventory skew is increased. The situation is reversed for the second tier (except of course around the point corresponding to zero inventory), since the skew is reduced through contributions from both the defensive and offensive sides, as the market maker aims to limit information leakage. The effect is particularly pronounced for top-of-book quotes. However, since skew sniffers still contribute significantly to risk reduction through their high $\Lambda_{0}$, the effect diminishes for larger sizes. A decrease in the trading intensity of the second tier leads to a sharp reduction of skew by the market maker across the entire pricing ladder for the second tier with skew sniffers, as illustrated in Figure 2.131313It is worth noting that for larger values of the risk aversion parameter $\gamma$, the market maker tends to reduce the skew shown to skew sniffers through a more important change in the quote on the defensive side than on the offensive side (figure not shown). This is the consequence of the model chosen where changes in price linked to price reading are proportional to skew.
5.3 Adverse Selection
Let us now consider the case of adverse selection only. We have
$D^{k}_{\pm}\widehat{f}(q)=\frac{\sum_{\begin{subarray}{c}m\in\mathcal{N}\\ j\in\mathcal{K}\end{subarray}}\alpha^{m,j}\Lambda_{0}^{m,j}(\beta^{m,j}-\kappa^{m,j})e^{\frac{\beta^{m,j}}{\kappa^{m,j}}}}{\sum_{\begin{subarray}{c}m\in\mathcal{N}\\ j\in\mathcal{K}\end{subarray}}\Delta^{j}\Lambda_{0}^{m,j}\kappa^{m,j}}\left(\pm q+\frac{\Delta^{k}}{2}\right),$
and therefore the quote adjustment is given by
$\displaystyle\frac{\widehat{\mathfrak{d}}^{n,k,b/a*}(q)-\widehat{\delta}^{n,k,b/a*}(q)}{\varepsilon}$ $\displaystyle=$ $\displaystyle\frac{\sum_{\begin{subarray}{c}m\in\mathcal{N}\\ j\in\mathcal{K}\end{subarray}}\alpha^{m,j}\Lambda_{0}^{m,j}(\beta^{m,j}-\kappa^{m,j})e^{\frac{\beta^{m,j}}{\kappa^{m,j}}}}{\sum_{\begin{subarray}{c}m\in\mathcal{N}\\ j\in\mathcal{K}\end{subarray}}\Delta^{j}\Lambda_{0}^{m,j}\kappa^{m,j}}\left(\pm q+\frac{\Delta^{k}}{2}\right)$
$\displaystyle\mp\frac{q\pm\Delta^{k}}{\Delta^{k}}\frac{\beta^{n,k}-\kappa^{n,k}}{\kappa^{n,k}}\zeta^{n,k}\left(\widehat{\delta}^{n,k,b/a*}(q)\right).$
As discussed above, the first term represents the global response to adverse selection risk, for trades of size $\Delta^{k}$, uniformly across all tiers, whereas the second term is specific to each tier. The magnitude and sign of these effects depend critically on the relative slopes of the exponential functions defining the intensity and price impact functions – that is, on the differences $\beta^{m,j}-\kappa^{m,j}$ appearing in the first sum, and on the tier-specific difference $\beta^{n,k}-\kappa^{n,k}$ in the second term.
As an illustration, we consider two tiers with equal intensities, as in the previous example. The first tier is assumed to be standard, with $\alpha^{1,k}=0$, while the second tier is endowed with a size-uniform informational advantage, characterized by $\alpha^{2,k}=0.1$, and slow signals, i.e., $\beta^{2,k}<\kappa^{2,k}$, here with $\frac{\beta^{2,k}}{\kappa^{2,k}}=0.9$.
Figure 3 shows that, in the presence of such adverse selection, the market maker behaves as if they were less risk averse when it comes to the quotes shown to the first tier. In particular, they tighten the bid-ask spread for the first tier when inventory is zero. By contrast, the spread quoted to the second (informed) tier is widened. Nevertheless, the market maker can still exploit the top-of-book prices shown to the informed tier to probe price direction. The limited effect on larger-size quotes is partly due to our modelling assumptions, but the underlying mechanism is general: small risk-reducing trades with informed clients cause limited losses while increasing the value of the remaining inventory.
A key takeaway from this analysis is that an informed tier subject to adverse selection should not necessarily be avoided unlike what initial intuition could suggest. Rather, it can be leveraged as a form of “signal subscription,” beneficial for the risk management of the entire franchise.
The picture is symmetrically reversed for adverse selection with sharp footprints, i.e., $\beta^{2,k}>\kappa^{2,k}$ (not shown). In this case, the market maker widens spreads for the first tier, acting as if they were more risk averse. Top-of-book prices for risk-decreasing trades with the informed tier become less aggressive: the market maker can no longer extract valuable information from such trades and must instead focus on protection.
Figure 3: Optimal bid quote adjustment, $\varepsilon^{-1}\left(\widehat{\mathfrak{d}}^{n,k,b*}(q)-\widehat{\delta}^{n,k,b*}(q)\right)$, for regular tier and tier with adverse selection. Risk aversion $\gamma=10^{-5}$ bp-1 M-1, other parameters are defined in the text.
Conclusion
In this article, we have proposed a tractable and implementable stochastic optimal control framework that enables market makers to adjust their quotes with greater awareness of informational risk arising from both price reading and adverse selection.
Using perturbation analysis, we have found that the impact of informational risk can be split into two components: (i) a global component related to induced changes of the value function that applies to pricing for all clients, and (ii) an individual component that applies to specific tiers to mitigate their perceived informational risk.
In the presence of price reading, the market maker acts as if they were more risk averse when showing prices to their global franchise. At the same time, they reduce the informational content of their quotes shown to skew sniffers. The extent of skew reduction strongly depends on liquidity, specifically on the contribution of price readers to market maker’s risk reduction. In particular, it is not beneficial to show skew to price readers with little liquidity contribution.
Optimal response to adverse selection depends on the relative shape of informed flow intensity and price impact functions. The market maker is able to become less risk averse for their global franchise when informed clients trade on long-term signals. Top-of-book prices for risk-reducing trades will be very attractive for informed clients. This is the cost the market maker is willing to pay to probe price direction for the benefit of a better risk management of their overall portfolio. On the contrary, adverse selection with sharp signals forces the market marker to be protective by showing defensive quotes even for risk-reducing trades with informed clients and behave as if they were more risk averse when it comes to their global franchise.
Acknowledgment
We would like to express our sincere gratitude to all those who made possible the research that led to this paper. The authors are especially grateful to Richard Anthony (HSBC).
Statement on Funding
The research conducted for this paper benefited from the financial support of the HSBC FX Research Initiative, a Research Initiative under the aegis of Institut Louis Bachelier, in partnership with HSBC UK. The views expressed are those of the authors and do not necessarily reflect the views or the practices at HSBC.
Statement on Data Availability
The data that support the findings of this article are available from the authors upon reasonable request.